# Parameter estimation for Gibbs distributions

### v9

@@ 5.2 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the log-concave setting
As we have mentioned, there is an alternative algorithm to estimate counts in the log-concave setting:
@@
3 set $w_{0}\leftarrow\min\{w_{0},\delta,1/n\}$ and set $w_{t}\leftarrow\min\{w_{t},\delta,1/n\}$
@@
7for $j=0,\dots,k_{1}$ do                 set $\hat{\pi}(j)=\hat{Q}(\beta_{\min})e^{-\beta_{\min}j}\hat{\mu}_{\beta_{\min}}(j)$
for $j=k_{t}+1,\dots,n$ do           set $\hat{\pi}(j)=\hat{Q}(\beta_{\max})e^{-\beta_{\max}j}\hat{\mu}_{\beta_{\max}}(j)$
@@
For correctness, suppose that $\mathcal{I}$ is proper and the call to PratioCoveringSchedule is good and every value $j\in\mathcal{H}$ is well-estimated by every iteration of Line 4. By specification of the subroutines, these events hold with probability at least $1-\gamma$. We claim that the preconditions of Lemma 14 then hold for all $j\in\mathcal{H}$. There are three cases to consider.
@@ 6.1 Constructing a preschedule
For Section 6, let us fix constants $\tau\in(0,\frac{1}{2})$, $\lambda>1$, and define parameter $\phi=\frac{\tau}{\lambda^{3}\rho}$. Thus, $\phi=\Theta(\frac{1}{\log n})$ in the general setting and $\phi=\Theta(1)$ in the log-concave setting. We introduce basic terminology and definitions.
@@
A segment is a tuple $\theta=(\beta,\sigma)$ where $\beta\in[\beta_{\min},\beta_{\max}]$, and $\sigma$ is an $\mathcal{H}$-interval. We say $\theta$ is $\phi$-proper (or just proper if $\phi$ is understood) if it satisfies the following two properties:
@@
We say that $\cal J$ is $\phi$-proper if all segments $\theta_{i}$ are $\phi$-proper. The main idea of the algorithm is to maintain a sequence of proper segments satisfying properties (I1) and (I2), and grow it until it satisfies (I0). Details are provided below:
@@
return $\mathcal{J}$     Subroutine ${\tt FindSegment}(\beta,[h_{{\tt left}},a_{{\tt left}}],[a_{{\tt right}},h_{{\tt right}}])$:
9 let $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\tfrac{1}{2}\log\lambda,\bar{p},\frac{1}{4(n+2)^{2}})$ where $\bar{p}=\frac{\phi}{h_{{\tt right}}-h_{{\tt left}}+1}$.
@@
return segment $\theta=(\beta,\sigma)$ where $\sigma=[k^{-},k^{+}]$
Algorithm 10 ${\tt FindSegment}(\beta,[h_{{\tt left}},a_{{\tt left}}],[a_{{\tt right}},h_{{\tt right}}])$.
@@
At each iteration, the uncovered set $[0,n]-\bigcup_{(\beta,\sigma)\in\mathcal{J}}\sigma$ is a union of disjoint open intervals (“gaps”). The interval $L=(\sigma^{+}_{{{\tt left}}},\sigma^{-}_{{{\tt right}}})$ at an iteration is one of these gaps. From the new segment produced by FindSegment, this gap $L$ gets replaced at the next iteration by two new gaps $L^{\prime},L^{\prime\prime}$. Overall, the set of gaps over all the iterations forms a laminar family of intervals. Between gaps, the “filled” regions have strictly increasing values of $\beta$. Thus at Line 4 we have $\beta_{\tt left}<\beta_{\tt right}$ strictly.
@@
By Theorem 11, the ${\tt Balance}$ subroutines have cost $O(n\log(nq))$. The two calls to FindSegment at Line 1 have cost $O(\rho n\log n)$. Let ${\cal J}_{i}$ be the sequence and $M_{i},\sigma_{{{\tt left}},i},\sigma_{{{\tt right}},i}$ be the variables at the $i^{\text{th}}$ iteration and let $A_{i}=[\sigma_{{{\tt left}},i}^{-},\sigma_{{{\tt right}},i}^{+}]$ and $L_{i}=(\sigma_{{\tt left},i}^{+},\sigma_{{\tt right},i}^{-})$. We can observe that the call to ${\tt FindSegment}$ at the $i^{\text{th}}$ iteration has cost $O(|A_{i}|\rho\log n)$, noting that $\sigma_{{\tt right},i}^{+}>\sigma_{{\tt left},i}^{-}$ strictly. We now show that $\sum_{i}|A_{i}|=O(n\log n)$, which will yield the claim about the complexity.
For each $k\in{\cal H}$ define $I^{-}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}<k\}$, and consider $i,j\in I^{-}_{k}$ with $i<j$. We claim that intervals $L_{i},L_{j}$ overlap. For, suppose $L_{j}$ comes strictly after $L_{i}$. Since $L_{j}$ is always selected from a gap at iteration $j$, it cannot intersect with $\sigma_{{\tt right},i}$, and in particular $\sigma_{{\tt left},j}^{+}\geq\sigma_{{\tt right},i}^{+}$. In this case, by definition of $M_{j}$ and $I_{k}$, we would have $k>M_{j}>\sigma_{{\tt left},j}^{+}$ and $k\leq\sigma_{{\tt right},i}^{+}$, a contradiction.
On the other hand, suppose $L_{j}$ comes strictly before $L_{i}$. Again, since $L_{j}$ is selected from a gap, it comes strictly before $\sigma_{{\tt left},i}$ and so $\sigma_{{\tt right},j}^{-}\leq\sigma_{{\tt left},i}^{-}$. By Property (I2) this implies $\sigma_{{\tt right},j}$ comes before $\sigma_{{\tt left},i}$ and so also $\sigma_{{\tt right},j}^{+}\leq\sigma_{{\tt left},i}^{+}$. Then by definition of $M_{i}$ and $I_{k}^{-}$ we would have $\sigma_{{\tt left},i}^{+}<M_{i}<k$ and $k\leq\sigma_{{\tt right},j}^{+}$, again a contradiction.
So $L_{i}$ and $L_{j}$ overlap. Since $M_{i}$ is chosen near the median of $L_{i}$, we must have $|L_{j}|\leq\tfrac{1}{2}|L_{i}|$. As this holds for all pairs $i,j\in I^{-}_{k}$, we conclude that $|I^{-}_{k}|\leq 1+\log_{2}n$. Similarly, $|I^{+}_{k}|\leq 1+\log_{2}n$ where $I^{+}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}>k\}$. It remains to observe that $\sum_{i}|A_{i}|=\sum_{k\in{\cal H}}|I^{-}_{k}\cup I^{+}_{k}|$. ∎
The main result for this theorem, which is technically involved, is the following:
@@
Indeed, we will show by induction on iteration count that the following two additional invariants are preserved in this case:
@@
Now we consider some call to FindSegment, made at Line 1 or some iteration of Line 7. By induction, properties (I3) and (I4) hold for all segments in $\mathcal{J}$ up to this point. We need to show that the segment $\theta$ returned by FindSegment also satisfies these properties. The cases when FindSegment is called in Line 1, or in Line 7 when $\beta\in\{\beta_{{\tt left}},\beta_{{\tt right}}\}$, are handled very differently from the main case, which is Line 7 with $\beta\in(\beta_{{\tt left}},\beta_{{\tt right}})$ strictly. In the former cases, there is no “free choice” for the left-margin $k^{-}$ or right-margin $k^{+}$ respectively; for instance, when $\beta=\beta_{{\tt left}}$, then $h_{{\tt left}}=a_{{\tt left}}=\sigma_{{\tt left}}^{-}$ and our only choice is to set $k^{-}=\sigma_{{\tt left}}^{-}$.
@@
$\mu_{\beta_{{\tt left}}}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta_{{\tt left}}}(h_{{\tt left}})$ (14)
@@
$\mu_{\beta}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta}(h_{{\tt left}})$
@@
$\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\leq\frac{a_{{\tt left}}+1-h_{{\tt left}}}{(a_{{\tt left}}+1-h_{{\tt left}})+(h_{{\tt left}}-i)}=\frac{a_{{\tt left}}-h_{{\tt left}}+1}{a_{{\tt left}}-i+1}\qed$
@@
(a) If the call is left-free, then $(a_{{\tt left}}-k^{-}+1)\cdot\hat{\mu}_{\beta}(k^{-})\geq\phi\sqrt{\lambda}$ and $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}$.
(b) If the call is right-free, then $(k^{+}-a_{{\tt right}}+1)\cdot\hat{\mu}_{\beta}(k^{+})\geq\phi\sqrt{\lambda}$ and $\mu_{\beta}(k^{+})\geq\hat{\mu}_{\beta}(k^{+})/\sqrt{\lambda}$.
@@
By Lemma 34, there is $k\in[h_{{\tt left}},a_{{\tt left}}]$ with $\mu_{\beta}(k)\geq\frac{\phi\lambda^{2}}{a_{{\tt left}}-k+1}\geq\bar{p}$. Since Line 10 well-estimates $k$, we have $\Phi(k)\geq\hat{\mu}_{\beta}(k)\geq\mu_{\beta}(k)/\sqrt{\lambda}\geq\frac{\phi\lambda^{3/2}}{a_{{\tt left}}+1-k}$. Since $k^{-}$ is chosen as the argmax, $(a_{{\tt left}}+1-k^{-})\Phi(k^{-})\geq(a_{{\tt left}}+1-k)\Phi(k)\geq\phi\lambda^{3/2}$. So $\hat{\mu}_{\beta}(k^{-})\geq\Phi(k^{-})/\lambda\geq\frac{\phi\sqrt{\lambda}}{a_{{\tt left}}+1-k^{-}}\geq\bar{p}$. Since Line 10 well-estimates $k^{-}$, this implies $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}$. ∎
@@
If the call is left-free, then Proposition 35 gives $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}\geq\frac{\phi}{a_{{\tt left}}+1-k^{-}}\geq\frac{\phi}{|\sigma|}$. ∎
@@
The segment $\theta=(\beta,\sigma)$ produced at the given iteration satisfies properties (I4).
@@
By Proposition 35, $(a_{{\tt left}}-k^{-}+1)\hat{\mu}_{\beta}(k^{-})\geq\phi\sqrt{\lambda}$. In particular, the RHS of Eq. (19) is at least $\bar{p}$. Since Line 1 well-estimates $i$ and since $\hat{\mu}_{\beta}(k^{-})\leq\sqrt{\lambda}\mu_{\beta}(k^{-})$, this implies
@@ 6.2 Converting the preschedule into a covering schedule
There are two steps to convert the preschedule into a covering schedule. First, we throw away redundant intervals; second, we “uncross” the adjacent intervals. While we are doing this, we also check if the resulting schedule is proper; if not, we will discard it and generate a new preschedule from scratch. Here $\nu>0$ is some arbitrary constant.
@@
3while $\mathcal{J}$ contains any index $i\in\{1,\dots,t-1\}$ with $\sigma_{i+1}^{-}\leq\sigma_{i-1}^{+}$ do
@@
5let $\mathcal{J}^{\prime}=((\beta_{0},\sigma_{0}),\dots,(\beta_{t},\sigma_{t}))$ be the preschedule after these modification steps. for $i=0,\dots,t$ do let $\hat{\mu}_{\beta_{i}}\leftarrow{\tt Sample}(\beta_{i};\tfrac{\nu}{2},e^{-\nu/2}w_{i},\frac{\gamma}{4(t+1)})$ where $w_{i}=\phi/|\sigma_{i}|$
@@
7       if  $\exists\;k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ with  $\hat{\mu}_{\beta_{i-1}}(k)\geq e^{-\nu/2}w_{i-1}$ and $\hat{\mu}_{\beta_{i}}(k)\geq e^{-\nu/2}w_{i}$  then
@@
return covering schedule $\mathcal{I}=(\beta_{0},e^{-\nu}w_{0},k_{1},\beta_{1},e^{-\nu}w_{1},k_{2},\dots,k_{t},\beta_{t},e^{-\nu}w_{t})$
@@
The algorithm TryToUncross satisfies the following properties:
(a) The output is either ERROR or a covering schedule $\mathcal{I}$ with ${\tt InvWt}(\mathcal{I})\leq\frac{2e^{\nu}(n+1)}{\phi}$.
(b) The output is an improper covering schedule with probability at most $\gamma$.
(c) If $\mathcal{J}$ is $\phi$-proper, then it outputs ERROR with probability at most $\gamma$.

### v11

@@ 5.2 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the log-concave setting
3
4set $w_{0}\leftarrow\min\{w_{0},\delta,1/n\}$ and set $w_{t}\leftarrow\min\{w_{t},\delta,1/n\}$
@@
9for $j=0,\dots,k_{1}$ do                 set $\hat{\pi}(j)=\hat{Q}(\beta_{\min})e^{-\beta_{\min}j}\hat{\mu}_{\beta_{\min}}(j)$
10 for $j=k_{t}+1,\dots,n$ do           set $\hat{\pi}(j)=\hat{Q}(\beta_{\max})e^{-\beta_{\max}j}\hat{\mu}_{\beta_{\max}}(j)$
@@
For correctness, suppose $\mathcal{I}$ is proper and the call to PratioCoveringSchedule is good and every value in $\mathcal{H}$ is well-estimated by every iteration of Line 4. By specification of the subroutines, these events hold with probability at least $1-\gamma$. We claim that the preconditions of Lemma 14 then hold for any $j\in\mathcal{H}$. There are three cases to consider.
@@ 6.1 Constructing a preschedule
We introduce basic terminology and definitions.
@@
A segment is a tuple $\theta=(\beta,\sigma)$ where $\beta\in[\beta_{\min},\beta_{\max}]$, and $\sigma$ is an $\mathcal{H}$-interval. We say $\theta$ is $\phi$-proper for parameter $\phi$ (or just proper if $\phi$ is understood) if it satisfies the following two properties:
@@
We say that $\cal J$ is $\phi$-proper if all segments $\theta_{i}$ are $\phi$-proper.
Let us fix constants $\tau\in(0,\frac{1}{2})$, $\lambda>1$, and define parameter
$\phi=\frac{\tau}{\lambda^{3}\rho}.$
Thus, $\phi=\Theta(\frac{1}{\log n})$ in the general setting and $\phi=\Theta(1)$ in the log-concave setting. The main idea of the algorithm is to maintain a sequence of proper segments satisfying properties (I1) and (I2), and grow it until it satisfies (I0). Details are provided below:
@@
9return $\mathcal{J}$
    Subroutine ${\tt FindSegment}(\beta,[h_{{\tt left}},a_{{\tt left}}],[a_{{\tt right}},h_{{\tt right}}])$:
10
11let $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\tfrac{1}{2}\log\lambda,\frac{\phi}{h_{{\tt right}}-h_{{\tt left}}+1},\frac{1}{4(n+2)^{2}})$.
@@
14 return segment $\theta=(\beta,\sigma)$ where $\sigma=[k^{-},k^{+}]$
@@
At each iteration, we say that $\mathcal{J}$ has a gap $(\sigma_{i}^{+},\sigma_{i+1}^{-})$ whenever $\sigma_{i}^{+}<\sigma_{i+1}^{-}$ strictly. The interval $G=(\sigma^{+}_{{{\tt left}}},\sigma^{-}_{{{\tt right}}})$ is one of these gaps. From the new segment produced by FindSegment, this gap $G$ gets replaced at the next iteration by two new gaps $G^{\prime},G^{\prime\prime}$. Overall, the set of gaps over all the iterations forms a laminar family of intervals. Between gaps, the “filled” regions have strictly increasing values of $\beta$. Thus at Line 4 we have $\beta_{\tt left}<\beta_{\tt right}$ strictly.
@@
By Theorem 10, the ${\tt Balance}$ subroutines have cost $O(n\log(nq))$. The two calls to FindSegment at Line 1 have cost $O(n\rho\log n)$. Let ${\cal J}_{i}$ be the sequence and $M_{i},\sigma_{{{\tt left}},i},\sigma_{{{\tt right}},i}$ be the variables at the $i^{\text{th}}$ iteration and let $A_{i}=[\sigma_{{{\tt left}},i}^{-},\sigma_{{{\tt right}},i}^{+}]$ and $G_{i}=(\sigma_{{\tt left},i}^{+},\sigma_{{\tt right},i}^{-})$. Observe that the call to ${\tt FindSegment}$ at the $i^{\text{th}}$ iteration has cost $O(|A_{i}|\rho\log n)$. We now show that $\sum_{i}|A_{i}|=O(n\log n)$, which will yield the claim about the complexity.
For each $k\in{\cal H}$ define $I^{-}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}<k\}$, and consider $i,j\in I^{-}_{k}$ with $i<j$. We claim that intervals $G_{i},G_{j}$ overlap. For, suppose $G_{j}$ comes strictly after $G_{i}$. Since $G_{j}$ is always selected from a gap at iteration $j$, it cannot intersect with $\sigma_{{\tt right},i}$, and in particular $\sigma_{{\tt left},j}^{+}\geq\sigma_{{\tt right},i}^{+}$. In this case, by definition of $M_{j}$ and $I_{k}$, we would have $k>M_{j}>\sigma_{{\tt left},j}^{+}$ and $k\leq\sigma_{{\tt right},i}^{+}$, a contradiction.
On the other hand, suppose $G_{j}$ comes strictly before $G_{i}$. Again, since $G_{j}$ is selected from a gap, it comes strictly before $\sigma_{{\tt left},i}$ and so $\sigma_{{\tt right},j}^{-}\leq\sigma_{{\tt left},i}^{-}$. By Property (I2) this implies $\sigma_{{\tt right},j}$ comes before $\sigma_{{\tt left},i}$ and so also $\sigma_{{\tt right},j}^{+}\leq\sigma_{{\tt left},i}^{+}$. Then by definition of $M_{i}$ and $I_{k}^{-}$ we would have $\sigma_{{\tt left},i}^{+}<M_{i}<k$ and $k\leq\sigma_{{\tt right},j}^{+}$, again a contradiction.
So $G_{i}$ and $G_{j}$ overlap. Since $M_{i}$ is chosen near the median of $G_{i}$, we must have $|G_{j}|\leq\tfrac{1}{2}|G_{i}|$. As this holds for all pairs $i,j\in I^{-}_{k}$, we conclude that $|I^{-}_{k}|\leq 1+\log_{2}n$. Similarly, $|I^{+}_{k}|\leq 1+\log_{2}n$ where $I^{+}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}>k\}$. It remains to observe that $\sum_{i}|A_{i}|=\sum_{k\in{\cal H}}|I^{-}_{k}\cup I^{+}_{k}|$. ∎
The final result in order to analyze Algorithm 10, which is technically involved, is the following:
@@
We will show by induction on iteration count that the following two additional invariants are preserved in this case:
@@
Now consider some call to FindSegment, either at Line 1 or Line 7. By induction, properties (I3) and (I4) hold for all segments in $\mathcal{J}$ up to this point. We must show that the segment $\theta$ returned by FindSegment also satisfies these properties. The cases when FindSegment is called in Line 1, or in Line 7 when $\beta\in\{\beta_{{\tt left}},\beta_{{\tt right}}\}$, are handled very differently from the main case, which is Line 7 with $\beta\in(\beta_{{\tt left}},\beta_{{\tt right}})$ strictly. In the former cases, there is no “free choice” for the left-margin $k^{-}$ or right-margin $k^{+}$ respectively; for instance, when $\beta=\beta_{{\tt left}}$, then $h_{{\tt left}}=a_{{\tt left}}=\sigma_{{\tt left}}^{-}$ and our only choice is to set $k^{-}=\sigma_{{\tt left}}^{-}$.
@@
$\mu_{\beta_{{\tt left}}}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta_{{\tt left}}}(h_{{\tt left}}).$ (14)
@@
$\mu_{\beta}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta}(h_{{\tt left}}).$
@@
$\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\leq\frac{a_{{\tt left}}+1-h_{{\tt left}}}{(a_{{\tt left}}+1-h_{{\tt left}})+(h_{{\tt left}}-i)}=\frac{a_{{\tt left}}-h_{{\tt left}}+1}{a_{{\tt left}}-i+1}.\qed$
@@
(a) If the call is left-free, then $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}\geq\frac{\phi}{a_{{\tt left}}-k^{-}+1}$.
(b) If the call is right-free, then $\mu_{\beta}(k^{+})\geq\hat{\mu}_{\beta}(k^{+})/\sqrt{\lambda}\geq\frac{\phi}{k^{+}-a_{{\tt right}}+1}$.
@@
Let $\bar{p}=\frac{\phi}{h_{{\tt right}}-h_{{\tt left}}+1}$. By Lemma 34, there is $k\in[h_{{\tt left}},a_{{\tt left}}]$ with $\mu_{\beta}(k)\geq\frac{\phi\lambda^{2}}{a_{{\tt left}}-k+1}\geq\bar{p}$. Since Line 10 well-estimates $k$, we have $\hat{\mu}_{\beta}(k)\geq\mu_{\beta}(k)/\sqrt{\lambda}\geq\frac{\phi\lambda^{3/2}}{a_{{\tt left}}+1-k}$. Since $k^{-}$ is chosen as the argmax, $(a_{{\tt left}}+1-k^{-})\Phi(k^{-})\geq(a_{{\tt left}}+1-k)\Phi(k)\geq\phi\lambda^{3/2}$. So $\hat{\mu}_{\beta}(k^{-})\geq\Phi(k^{-})/\lambda\geq\frac{\phi\sqrt{\lambda}}{a_{{\tt left}}+1-k^{-}}\geq\bar{p}$. Since Line 10 well-estimates $k^{-}$, this implies $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}$. ∎
@@
If the call is left-free, then Proposition 35 gives $\mu_{\beta}(k^{-})\geq\frac{\phi}{a_{{\tt left}}-k^{-}+1}\geq\frac{\phi}{|\sigma|}$. ∎
@@
The segment $\theta=(\beta,\sigma)$ produced at the given iteration satisfies property (I4).
@@
By Proposition 35, $(a_{{\tt left}}-k^{-}+1)\hat{\mu}_{\beta}(k^{-})\geq\phi\sqrt{\lambda}$. In particular, the RHS of Eq. (19) is at least $\bar{p}=\frac{\phi}{h_{{\tt right}}-h_{{\tt left}}+1}$. Since Line 1 well-estimates $i$ and since $\hat{\mu}_{\beta}(k^{-})\leq\sqrt{\lambda}\mu_{\beta}(k^{-})$, this implies
@@ 6.2 Converting the preschedule into a covering schedule
There are two steps to convert the preschedule into a covering schedule. First, we throw away redundant intervals; second, we “uncross” the adjacent intervals. While we are doing this, we also check if the resulting schedule is proper; if not, we will discard it and generate a new preschedule from scratch.
@@
3
4while $\mathcal{J}$ contains any index $i\in\{1,\dots,t-1\}$ with $\sigma_{i+1}^{-}\leq\sigma_{i-1}^{+}$ do
@@
6
7let $\mathcal{J}^{\prime}=((\beta_{0},\sigma_{0}),\dots,(\beta_{t},\sigma_{t}))$ be the preschedule after these modification steps.
8for $i=0,\dots,t$ do let $\hat{\mu}_{\beta_{i}}\leftarrow{\tt Sample}(\beta_{i};\tfrac{1}{2}\log\lambda,\frac{w_{i}}{\lambda},\frac{\gamma}{4(t+1)})$ where $w_{i}=\phi/|\sigma_{i}|$
@@
10       if  $\exists\;k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ with  $\hat{\mu}_{\beta_{i-1}}(k)\geq w_{i-1}/\sqrt{\lambda}$ and $\hat{\mu}_{\beta_{i}}(k)\geq w_{i}/\sqrt{\lambda}$  then
@@
return covering schedule $\mathcal{I}=(\beta_{0},w_{0}/\lambda,k_{1},\beta_{1},w_{1}/\lambda,k_{2},\dots,k_{t},\beta_{t},w_{t}/\lambda)$
@@
The algorithm ${\tt TryToUncross}(\mathcal{J},\gamma)$ satisfies the following properties:
(a) The output is either ERROR or a covering schedule $\mathcal{I}$ with ${\tt InvWt}(\mathcal{I})\leq\frac{2\lambda(n+1)}{\phi}$.
(b) Irrespective of $\mathcal{J}$, it outputs an improper covering schedule with probability at most $\gamma$.
(c) If $\mathcal{J}$ is $\phi$-proper, it outputs ERROR with probability at most $\gamma$.